(1 pt) Table 12-2 of Harris gives formation constants (as log K) for many metal ions. Calculate conditional formation constants (Kf) for EDTA complexation of Mn2+ and Cu2+ at pH 6.00. Explain why we can no longer titrate Mn2+ but can still titrate Cu2+ if we reduce the pH to 4.00.
TABLE 12-2 Formation constants for metal-EDTA complexes Ion log Kr Ion log K Ion log Kr 2.95 V3+ 1.86 Cr3+ 0.8Mn3+ 9.7 Fe3+ 8.79 Co3 10.65Zr*+ 8.72 | Нґ+ 7.88 Vo2+ 7.4 23.1Ag 18,08 TI 15.36 Pd2+ 12.7 Zn2+ 13.6 Cd2+ 13.89 Hg2 4.30 Sn2+ 16.45Pb2+ 18.4 18.78 Ga 21.3 In3+ 25.9 TI+ 23.4Bi3 25.2 Ce+ 25. Pr3 41.4Nd3+ 29.3Pm3 29.5 Sm3+ 8.7 Eu3+ 15.5 Gd+ 35.3 27.8 15.93 16.30 16.51 16.9 17.06 17.25 17.35 17.87 18.30 18.56 18.89 19.32 19.49 19.74 23.2 25.7 Mg2* 2+ 7.20 Tb3 6.41 Dy+ 5.6Ho3+ 6.5 Er 16.5 Tm3+ 21.5 Yb3+ 18.3° | Lu3+ 18.0 Th 16.4 U4+ 21.7 24.9 2+ 2+ Co2+ Al 2+ 3+ NOTE: The stability constant is the equilibrium constant for the reaction M"Y4MY"4. Values in table apply at 25°C and ionic strength 0.1 M unless otherwise indicated a. 20°C, ionic strength- 0.1 M SOURCE: A. E. Martell, R. M. Smith, and R. J. Motekaitis, NIST Critically Selected Stability Constants of Metal Complexes, NIST Standard Reference Database 46, Gaithersburg, MD, 2001. b. 20°C, ionic strength-1 M
Answer #1

At pH 6, a​​​​​​ y​​​​​​ 4- = 2× 10-5

For Cu2+, LogKf = 18.78, Kf = 1018.78 = 6.0256 × 1018

Condition formation constant Kf' = a y​​​​​​4+ Kf = 1.205 × 1014

For Mn2+ LogKf = 13.89, Kf = 7.67 × 10^13

Condition formation constant Kf' = a​​​​​​ y​​​​​​4- Kf = 1.552 × 109

One general trend you should see is that as pH gets lower so does the K. At what point does it get so low that we can no longer do the titration?? As a general rule of thumb we use 108 as a cutoff. Thus K’ must be >=108 for an EDTA titration to work.

At pH 4, a y 4- = 4 × 10-9.

Kf' for Cu2+ = 4× 10-9 × 6.0256 × 1018  = 2.4102 × 1010

Kf' for Mn2+ = 4× 10-9 × 7.67 × 1013 = 3.105 × 105

So at pH 4, Kf' for Mn2+ is less than 108 , we can't do the titration of Mn2+ while Kf' for Cu2+ is greater than 108 , so we can do the titration of Cu2+ at pH 4.